The average height of 12 children is 124.7 with a standard deviation of 7.5. Assuming

that the heights are normally distributed, what percent of the students have a height of

132.2 and longer?


1
Expert's answer
2022-05-02T13:51:34-0400
"P(X\ge132.2)=1-P(Z<\dfrac{132.2-124.7}{7.5})"

"=1-P(Z<1)\approx0.158655"

15.8655%